# Tora: Trajectory-oriented Diffusion Transformer for Video Generation

### v1

@@ Introduction
Quantitative and Qualitative Results
We compare our method against prevalent motion-guided video generation approaches. The evaluation is conducted under three settings: 16, 64, and 128 frames, all with a resolution of 512 $\times$ 512 for fair comparison. The provided trajectories are clipped to accommodate different evaluated video lengths. For most U-Net based methods, we employ sequenced inference, wherein the last frame generated by the previous batch serves as the visual condition for the current batch, aligning with their inference settings. As illustrated in Table 1, under the 16-frame setting commonly employed by U-Net based methods, MotionCtrl and DragNUWA exhibit better alignment with the provided trajectories but are still weaker compared to our proposed Tora. As the frame count increases, the U-Net based methods show significant deviations in certain frames, with misalignment errors propagating and leading to deformation, motion blur, or object disappearance in subsequent sequences. In contrast, Tora demonstrates high robustness to varying frame numbers due to the integration of the transformer’s scaling ability. The movement generated by Tora is smoother and more aligned with the physical world. When evaluated under the 128-frame test setting, Tora’s trajectory accuracy surpasses other methods by a factor of 3 to 5, demonstrating its exceptional motion control capabilities. In Figure 4, we provide an analysis of the Trajectory Error across varying resolutions and durations. Unlike U-Net based models, which exhibit significant trajectory errors over time, Tora shows only a gradual increase in Trajectory Error as duration increases. This gradual increase in error aligns with the decrease in video quality observed in the DiT model as the duration extends. The results clearly indicate that our method maintains effective trajectory control over longer durations.
Figure 5: Qualitative Comparisons on Trajectory Control. All methods are capable of generating objects that follow the given trajectory. However, Tora not only adheres precisely to the specified trajectory but also produces smoother movement that conforms to the physics world. Please visit our supplemental materials for detailed video results.
Method FVD CLIPSIM TrajError
@@
Table 2: Evaluations of the impact of different trajectory compression methods. The 3D VAE yields superior results in motion-conditioned encoding.
Figure 5 presents a comparative analysis of our proposed method against mainstream motion control techniques. In the first scenario, which involves the co-movement of two men, all methods are able to generate relatively accurate motion trajectories. However, our approach demonstrates superior visual quality. This advantage is primarily due to the utilization of longer sequence frames, which facilitate smoother motion trajectories and more realistic background rendering. For instance, in our generated bicycle scenario, the human legs exhibit realistic pedaling motions, whereas the output from DragNUWA features legs that appear almost horizontally floating, violating physical realism. Additionally, both DragNUWA and MotionCtrl suffer from significant motion blur towards the end of their videos. Furthermore, MotionCtrl introduces unintended camera movements during the riding sequence, despite the absence of camera movement conditions. In another case, DragNUWA shows severe deformation of the lantern as the provided trajectory rises and falls continuously. Although MotionCtrl’s trajectory is relatively accurate, the resulting video fails to match the expected depiction of two lanterns. Overall, our method not only adheres closely to the provided trajectories but also minimizes object deformation, thereby ensuring higher fidelity in motion representation.
Ablation study
In this section, we conduct serval ablation studies to analyze the effects of our design choices. All models are evaluated using 480p resolution, a 16:9 aspect ratio, and 204 frames.
Method FVD CLIPSIM TrajError
@@
Trajectory Compression. To incorporate the trajectory vector within the same latent space as video patches, we investigate three distinct methods for trajectory compression, as summarized in Table 2. The first method samples the mid-frame as the keyframe for successive 4-frame intervals and employs patch-unshuffle for spatial compression. Despite its simplicity, this method proves suboptimal for motion control, primarily due to the potential flow estimation errors when encountering rapid motions or occlusions. Additionally, the magnified dissimilarity between patches induced by the chosen frame interval exacerbates learning challenges. The second method utilizes average pooling to gather successive frames. While this captures a general sense of movement, it inadvertently sacrifices precision by homogenizing the trajectory’s direction and magnitude, thereby diluting critical motion details. To preserve the trajectory information between consecutive frames as much as possible, we further employ a 3D VAE to extract the global context of successive trajectory intervals. The trajectory data is visually translated into an RGB image format to leverage existing 3D VAE weights. Extensive training on a large scale of trajectory videos with this setup yields the most favorable outcomes, underscoring the efficacy of our tailored 3D VAE approach in trajectory compression.
Block design and integrated position of MGF. We train the three variant MFG blocks as previously described, with the results presented in Table 3. Notably, the adaptive norm block achieves lower FVD and Trajectory Error compared to both cross-attention and extra channel conditioning methods, while also exhibiting the highest computational efficiency. This superiority is attributed to its capability for dynamic feature adaptation based on diverse conditions without requiring strict alignment, a limitation often encountered with cross-attention. Moreover, it ensures temporal consistency by modulating the conditional information over time, which is crucial for injecting motion cues. In contrast, channel concatenation may cause information congestion, rendering motion signals less effective. During training, we observe that initializing the normalization layer as the identity function is important for achieving optimal performance. Additionally, we evaluate the integration position of the MGF module within the Spatial DiT and Temporal DiT blocks. Our findings reveal that embedding the MGF within the Temporal DiT block notably enhances trajectory motion control, as evidenced by a substantial decrease in Trajectory Error (from 23.39 to 14.25). This approach enhances the MGF’s efficiency in interacting with temporal dynamics, thereby significantly improving the fidelity of motion synthesis.
Training Strategies. We evaluate the effectiveness of the two-stage training approach, with the results summarized in Table 4. Training exclusively with dense optical flow proves ineffective as it fails to accurately capture the intricate details of the provided sparse trajectories. On the other hand, training solely with sparse trajectories offers limited information, making the learning process more challenging. By initially training with dense flows and subsequently fine-tuning with sparse trajectories, our model demonstrates improved adaptability to various types of trajectory data. This method not only enhances overall performance but also increases the model’s versatility in handling diverse motion patterns.
Motion-guidance FVD CLIPSIM TrajError
@@
Hybird 513 0.2358 14.25
Table 4: Ablation of the type of training trajectories. “Hybrid” denotes the proposed two-stage training strategy.
@@ Conclusion
This paper introduces Tora, the first trajectory-oriented Diffusion Transformer framework for video generation that integrates text, image, and trajectory conditions. Tora effectively encodes arbitrary trajectories into spacetime motion patches, which align with the scaling properties of DiT, thereby enabling more realistic simulations of physical world movements. By employing a two-stage training process, Tora achieves motion-controllable video generation across a wide range of durations, aspect ratios, and resolutions. Remarkably, it can generate high-quality videos that adhere to specified trajectories, producing up to 204 frames at 720p resolution. This capability underscores Tora’s versatility and robustness in handling diverse motion patterns while maintaining high visual fidelity. We hope our work provides a strong baseline for future research in motion-guided Diffusion Transformer methods.

### v3

@@ Results
Ablation study
We conduct several ablation studies to analyze the effects of our design choices. All models are evaluated using 480p resolution, a 16:9 aspect ratio, and 204 frames.
Trajectory Compression. To integrate the trajectory vector into the same latent space as video patches, we explore three different methods for trajectory compression, as summarized in Table 2. The first method samples the mid-frame as a keyframe for successive 4-frame intervals and uses patch-unshuffle for spatial compression. While simple, this approach is sub-optimal for motion control due to potential flow estimation errors during rapid movements or occlusions, and the increased dissimilarity between patches complicates learning. The second method employs average pooling to gather information from successive frames. Although this captures a general sense of movement, it sacrifices precision by averaging the trajectory’s direction and magnitude, diluting important motion details. To better preserve trajectory information between consecutive frames, we utilize a 3D VAE to extract the global context of successive trajectory intervals. The trajectory data is converted into RGB images to leverage existing 3D VAE weights. Extensive training on a large dataset of trajectory videos with this method yields the best results, highlighting the effectiveness of our customized 3D VAE approach for trajectory compression.
Method FVD ($\downarrow$) CLIPSIM ($\uparrow$) TrajError ($\downarrow$)
@@
Table 2: Evaluations of the impact of different trajectory compression methods.
Block design and integrated position of MGF. We train the three variant MGF blocks as previously described, with the results presented in Table 3. Notably, the adaptive norm block achieves lowest FVD and Trajectory Error, while also exhibiting the highest computational efficiency. This advantage stems from its ability to dynamically adapt features based on varying conditions without needing strict alignment, a common challenge with cross-attention. Additionally, it maintains temporal consistency by modulating conditional information over time, which is essential for incorporating motion cues. In contrast, channel concatenation can lead to information congestion, making motion signals less effective. We find that initializing the normalization layer as the identity function is vital for optimal performance. Additionally, placing the MGF module within the Temporal DiT block significantly enhances trajectory motion control, evidenced by a drop in Trajectory Error from 23.39 to 14.25.
Method FVD ($\downarrow$) CLIPSIM ($\uparrow$) TrajError ($\downarrow$)
@@
Training Strategies. We evaluate the two-stage training approach, with results in Table 4. Training only with dense optical flow is ineffective, as it fails to capture the details of sparse trajectories. Conversely, using only sparse trajectories provides limited information, complicating the learning process. By first training with dense flows and then fine-tuning with sparse trajectories, our model demonstrates better adaptability and versatility in managing various motion patterns, leading to improved overall performance.
Motion-guidance FVD ($\downarrow$) CLIPSIM ($\uparrow$) TrajError ($\downarrow$)
@@
Hybrid 513 0.2358 14.25
Table 4: Ablation of the type of training trajectories. “Hybrid” denotes the two-stage training strategy.
Figure 5: Qualitative Comparisons on Trajectory Control. All methods are capable of generating objects that follow the given trajectory. However, Tora not only adheres precisely to the specified trajectory but also produces smoother movement that conforms to the physics world.
@@ Conclusion
This paper introduces Tora, the first trajectory-oriented Diffusion Transformer framework for video generation. Tora effectively encodes arbitrary trajectories into spacetime motion patches, which align with the scaling properties of DiT, thereby enabling more realistic simulations of physical world movements. By employing a two-stage training process, Tora achieves motion-controllable video generation across a wide range of durations, aspect ratios, and resolutions. Remarkably, it can generate high-quality videos that adhere to specified trajectories, producing up to 204 frames at 720p resolution. This capability underscores Tora’s versatility and robustness in handling diverse motion patterns while maintaining high visual fidelity. We hope our work provides a strong baseline for future research in motion-guided Diffusion Transformer methods.
Appendix
This supplementary material offers additional results, comprehensive dataset information, and thorough analyses that bolster the findings and conclusions outlined in the main text. It is organized as follows:
  • Additional qualitative results.
  • Data pre-processing method.
  • Dataset details, regarding total quantity, total durations, etc.
  • Prompt refinement method.
  • Motion VAE training.
Qualitative Comparisons
While the main text focuses on quantitative comparisons with the motion-controllable video generation models and ablation studies on different designs for TE and MGF, here we provide further visual comparisons.
Compare with motion-controllable methods
Figure 6: Qualitative comparison between Tora and OpenSora. All results are generated under the same text and image conditions. Tora employs an appropriate trajectory that simulates real-world physics, leading to more coherent and stable motion.
Figure 5 provides a comparative analysis of our proposed method against mainstream motion control techniques. In the first scenario involving the coordinated movement of two individuals, all methods manage to generate motion trajectories that are relatively accurate. However, our approach stands out for its superior visual quality. This advantage is largely attributed to the use of longer sequence frames, which contribute to smoother motion trajectories and more realistic background rendering. For example, in our generated bicycle scenario, the human legs exhibit lifelike pedaling motions, while the output from DragNUWA shows legs that appear to float almost horizontally, compromising physical realism. Moreover, both DragNUWA and MotionCtrl encounter significant motion blur towards the end of their videos. Additionally, MotionCtrl introduces unintended camera movements during the riding sequence, despite the absence of any intended camera movement conditions. In another instance, DragNUWA suffers from severe deformation of the lantern as the provided trajectory oscillates up and down. Although MotionCtrl’s trajectory is relatively accurate, the resulting video does not align with the expected portrayal of two lanterns. Overall, our method not only adheres closely to the provided trajectories but also minimizes object deformation, thereby ensuring higher fidelity in motion representation.
Compare with OpenSora
Despite OpenSora’s impressive accomplishments, it faces challenges when creating long videos featuring complex motions, such as simultaneous movement of multiple objects, swinging, or circling. This often leads to incoherent or distorted foreground objects, negatively impacting visual quality. To our delight, we discovered that incorporating appropriate trajectory control into the DiT model offers a more effective constraining signal. This improvement markedly enhances video fluidity and preserves object fidelity, as demonstrated in Figure  6.
Figure 7: Generated videos employing different trajectory compression methods: (a) Sampling Keyframe; (b) Average Pooling; (c) 3D VAE.
In scenarios where a teddy bear is oscillating side to side on a skateboard or a rose is swirling in circular motions, OpenSora, which relies solely on textual directives for motion control, exhibits noticeable object deformations. In contrast, Tora excels at maintaining the inherent shape of the objects. Additionally, when managing the motion of multiple entities, such as a pair of jellyfish—one moving upward while the other moves downward, OpenSora demonstrates noticeable flickering, underscoring its limitations in handling complex movements. In conclusion, the integration of Tora’s motion signaling mechanism enhances both the controllability and stability of the synthesized videos.
Comparison of Different Trajectory Compression Methods
We train our proposed trajectory extractor using the various trajectory compression methodologies previously discussed. The comparisons of these methods are visually illustrated in Figure 7.
(a) Histogram of Caption Lengths.
(b) Histogram of Video Durations.
(c) The training batch size of every bucket (resolution, duration).
Figure 8: Overview of the training data distributions and batch sizes.
In key-frame sampling, while it successfully captures essential motion, it frequently leads to misalignment between video patches and motion patches, especially during rapid motion sequences. This misalignment hinders the generated objects from accurately tracking their trajectories, negatively impacting visual fluidity and overall quality. On the other hand, average pooling smooths out minor variations, resulting in a more consistent motion representation. However, in complex trajectories, such as S-shaped turns where consecutive frame directions are inconsistent, this approach may introduce artifacts because the physical relevance of optical flow decreases. In contrast, our proposed 3D VAE approach effectively compresses trajectory information into the video’s latent space. By training the 3D VAE on the large dataset with flow annotations, it successfully extracts the most relevant motion features for guidance, preserving the movement of successive frames to a significant extent. As evidenced in the results, this method significantly enhances the fluidity and coherence of the generated movements, producing visually compelling sequences that closely resemble natural motion.
Data Pre-processing
During the processing of the video datasets, constructing a high-quality training set is crucial as it significantly impacts the quality of the generated videos. The following is a detailed description of our data processing workflow, which includes steps such as invalid videos removal, resolution filtering, camera motion filtering, and assessing the degree of object motion.
Initially, during the dataset preparation phase, we remove invalid videos. This step aims to identify and discard videos that do not meet our established criteria, including those with encoding errors, a duration of zero, or low quality. We identify encoding errors and zero-duration videos by directly decoding them. Furthermore, we predict both the aesthetic score [^4] https://github.com/christophschuhmann/improved-aesthetic-predictor and the optical flow score (Xu et al. 2023) for each video. A video is deemed valid only if its aesthetic score exceeds 5.5 and its flow score is greater than 2.
Next, we perform resolution filtering. To ensure the effectiveness of subsequent study, we establish a minimum resolution standard of 720p. By checking the resolution of each video, we can eliminate those that fall below this threshold, thereby ensuring that the videos in our dataset possess adequate clarity and detail.
Subsequently, we perform camera motion filtering using a camera motion detector [^5] https://github.com/antiboredom/camera-motion-detector and a motion segmentor  (Zhao et al. 2022) to filter out videos with significant camera movement, which may distort the model’s ability to focus on the motion of the primary subjects. More specifically, the zoom detection threshold is set between 0.4 and 0.6. The detected camera movement angles, calculated based on the background from the motion segmentation results, are valid as follows:$[0^{\circ},20^{\circ}],[160^{\circ},200^{\circ}],[340^{\circ},360^{\circ}]$.
Finally, we analyze the magnitude score of the optical flow within the foreground, excluding those scenes that are mostly static or exhibit minimal movement. Moreover, dramatic object motions in some videos can cause significant optical flow deviations, interfering with trajectory training. Consequently, we retain these videos with a probability of $(1-score/100)$.
Through these rigorous filtering and processing steps, we successfully construct a high-quality video dataset suitable for subsequent training, providing a solid foundation for our study.
Dataset Details
This section offers an overview of the dataset used in this study, covering its origin and composition. We employ histograms and descriptive statistics to illustrate the dataset’s structure and distribution.
Training Data
The video data is sourced from the Panda-70M subset, Mixart, and internal videos. We initially collect 2.6M videos and apply the data preprocessing pipeline to filter the content, resulting in 631k eligible videos for training. An overview of the training dataset is presented in Table 5, which details the durations, resolution and FPS.
# Videos Clips 631053
Total Durations (hours) 2952.93
Average Shorter Edge Length 965.11
Average FPS 29.23
Table 5: Statistical information about the training data.
Additionally, Table 6 summarizes the mean and standard deviation for the durations, number of frames, and caption lengths. We also present histogram to show the distribution of the caption lengths and the durations of all video clips, as shown in the Figure 8(a) and Figure 8(b).
mean std
Durations (seconds) 16.85 19.58
#Frames 506.22 644.38
Caption Length (#word) 125.52 24.22
Table 6: Statistics of training set, regarding durations, number of frames, and caption lengths.
Figure 9: Visualization of Our Evaluation Dataset, highlighting 0%, 20%, 40%, 60%, 80%, and 100% of the total duration. Each center point of the annotated object masks is treated as a trajectory point. The number of trajectories in the tested video matches the number of annotated objects.
Drawing inspiration from OpenSora, we employ a multi-scale and mixed-duration training strategy, which involves training videos of various resolutions and lengths together. Specifically,we establish predefined buckets, each defined by a unique combination of (video resolution, duration). Videos are then assigned to the appropriate bucket according to their specific attributes. Note that videos of any aspect ratio will fall into these buckets if their total pixel count is within the specified statistical intervals. The parameter settings for the buckets adhere to the principle that lower resolutions correspond to longer durations, enabling Tora to adapt to videos of varying lengths. Notably, our preprocessing steps ensure that the shorter edge of each training video exceeds 720 pixels. To enable training across various scales, we shuffle the dataset and randomly select videos for downsampling to lower resolutions. Additionally, we employ different batch sizes for each bucket to balance the GPU load. The details of the buckets are presented in Figure 8(c).
Evaluation Data
Our evaluation dataset is primarily sourced from video object segmentation datasets (Xu et al. 2021; Qi et al. 2022; Pont-Tuset et al. 2017), which offer robust object motion critical for our analysis. To enhance the quality of our evaluation, we implement a camera motion filtering technique to select videos where the camera remains predominantly stable. This filtering process allows us to concentrate on where object motion is distinctly pronounced, thereby improving the reliability of our assessments. For each frame, we utilize the center of the annotated object masks as trajectory points, providing precise references for evaluating motion dynamics. Figure 9 presents several examples from our evaluation dataset, highlighting the diversity and relevance of the selected video sequences.
Prompt Refinement
We encourage users to provide detailed text prompts to achieve satisfactory video results. To ensure consistency in the distribution of text prompts during both training and testing phases, we utilize GPT-4o to refine simple testing prompts. The process of learning refined prompts for GPT-4o involves two key components. The first component is the task description, which clearly outlines the objectives for the model in generating enhanced content:
You need to refine user’s input prompt. The user’s input prompt is used for video generation task. You need to refine the user’s prompt to make it more suitable for the task. Here are some examples of refined prompts: $\downarrow$
a close-up shot of a woman applying makeup. she is using a black brush to apply a dark powder to her face. the woman has blonde hair and is wearing a black top. the background is black, which contrasts with her skin tone and the makeup. the focus is on her face and the brush, with the rest of her body and the background being out of focus. the lighting is soft and even, highlighting the texture of the makeup and the woman’s skin. there are no texts or other objects in the video. the woman’s expression is neutral, and she is looking directly at the camera. the video does not contain any action, as it is a still shot of a woman applying makeup. the relative position of the woman and the brush is such that the brush is in her hand and is being used to apply the makeup to her face. the video does not contain any other objects or actions. the woman is the only person in the video, and she is the main subject. the video does not contain any sound. the description is based on the visible content of the video and does not include any assumptions or interpretations. $\downarrow$
a professional setting where a woman is presenting a slide from a presentation. she is standing in front of a projector screen, which displays a bar chart. the chart is colorful, with bars of different heights, indicating some sort of data comparison. the woman is holding a pointer, which she uses to highlight specific parts of the chart. she is dressed in a white blouse and black pants, and her hair is styled in a bun. the room has a modern design, with a sleek black floor and a white ceiling. the lighting is bright, illuminating the woman and the projector screen. the focus of the image is on the woman and the projector screen, with the background being out of focus. there are no texts visible in the image. the relative positions of the objects suggest that the woman is the main subject of the image, and the projector screen is the object of her attention. the image does not provide any information about the content of the presentation or the context of the meeting. $\downarrow$
a serene scene in a park. the sun is shining brightly, casting a warm glow on the lush green trees and the grassy field. the camera is positioned low, looking up at the towering trees, which are the main focus of the image. the trees are dense and full of leaves, creating a canopy of green that fills the frame. the sunlight filters through the leaves, creating a beautiful pattern of light and shadow on the ground. the overall atmosphere of the video is peaceful and tranquil, evoking a sense of calm and relaxation. $\downarrow$
a moment in a movie theater. a couple is seated in the middle of the theater, engrossed in the movie they are watching. the man is dressed in a casual outfit, complete with a pair of sunglasses, while the woman is wearing a cozy sweater. they are seated on a red theater seat, which stands out against the dark surroundings. the theater itself is dimly lit, with the screen displaying the movie they are watching. the couple appears to be enjoying the movie, their attention completely absorbed by the on-screen action. the theater is mostly empty, with only a few other seats visible in the background. the video does not contain any text or additional objects. the relative positions of the objects are such that the couple is in the foreground, while the screen and the other seats are in the background. the focus of the video is clearly on the couple and their shared experience of watching a movie in a theater. $\downarrow$
a scene where a person is examining a dog. the person is wearing a blue shirt with the word "volunteer" printed on it. the dog is lying on its side, and the person is using a stethoscope to listen to the dog’s heartbeat. the dog appears to be a golden retriever and is looking directly at the camera. the background is blurred, but it seems to be an indoor setting with a white wall. the person’s focus is on the dog, and they seem to be checking its health. the dog’s expression is calm, and it seems to be comfortable with the person’s touch. the overall atmosphere of the video is calm and professional. $\downarrow$
The refined prompt should pay attention to all objects in the video. The description should be useful for AI to re-generate the video. The description should be no more than six sentences. The refined prompt should be in English.
Following that, GPT-4o is supplied with the testing captions for processing. This allows it to refine the prompts based on the initial task description, ensuring that the provided captions are more detailed and aligned with our objectives:
Generate the refined prompts for following inputs: $\downarrow$
A man rides on a huge fish, flying from the water into the sky. $\downarrow$
Two Jedi cats are fighting with each other in the forest. $\downarrow$
A polar bear with a black hat is walking on the Great Wall. $\downarrow$
A woman and a golden retriever are playing on the beach at sunset. $\downarrow$
Two roses sway together before a snow-covered mountain range.
Motion VAE Training
Given the absence of pre-existing networks tailored for video optical flow compression, training such a network from scratch presents significant challenges. Directly transferring the motion vectors to the image domain and applying a pretrained 3D VAE may hinder the model’s ability to effectively encode motion features, primarily due to domain discrepancies. To overcome this issue, we refine a motion-specific 3D VAE that is initialized from a pretrained model. Specifically, our motion 3D VAE is specifically initialized using the architecture of OpenSora’s VAE, which adapts the structure of Magvit-v2. This VAE has a substantial parameter count of 384 million, effectively leveraging the capabilities of a well-established network. Our training data is sourced from a combination of datasets annotated with optical flow information (Mehl et al. 2023; Mayer et al. 2016; Ranjan et al. 2020; Cabon, Murray, and Humenberger 2020). We fine-tune the motion 3D VAE for 200,000 iterations with a batch size of 1. The training video size is set to a random number of frames, capped at 34. This setting aligns with the OpenSora video VAE, improving compatibility between the motion VAE and the video VAE and ensuring a cohesive training process. We utilize PSNR, SSIM and trajectory error to evaluate reconstruction quality and motion-controllable ability. The performance differences between the pure video VAE and our fine-tuned model are presented in Table 7.
[…]
Pure Video VAE 27.34 0.842 17.09
Our VAE 28.76 0.860 14.25
Table 7: The performance comparison of different 3D VAE.